Following is partial information for Delamunte Industries for the month of August: Work In Process Balance, August 1 Direct materials used Direct labor Manufacturing overhead applied $ 128,000 122,800 166,000 151,000 Jobs finished during August are summarized here: Cost of Jobs Job # 234 237 231 246 Completed $71,950 32,800 60,000 27,000 Required At the end of August, only one job, Job 248, was still in process. The direct labor cost incurred on Job 248 as of August 31 was $23,500 1. Calculate the predetermined overhead rate that was used to apply manufacturing overhead to jobs during August. Round your answer to 2 decimal places.) Predetermined Overhead Rate per Direct Labor dollar 2. Determine the balance in Work in Process at the end of August. k in Process 3. Determine how much of the direct materials cost and applied overhead would be recorded for Job 248 as of August 31. (Round your Predetermined Overhead Rate to 2 decimal places.) Direct Materials Cost Applied Overhead
Answer #1

1. Predetermined overhead rate per direct labor dollar = total manufacturing overhead applied / total direct labor cost

= 151000/166000 = $0.91 per direct labor dollar

Predetermined overhead rate $0.91 Per direct labor dollar

2.

Work in progress at the end of August (see calculation in notes ) $376050

Notes :-

1 Balance in work in progress as on Aug 31 will be calculated as follows :-

Balance WIP as on aug 1 128000
Direct material used 122800
Direct labor 166000
Manufacturing overhead applied 151000
Total cost on all jobs till the end of aug $567800
(-) cost of following fininshed jobs
- job 234 (71950)
- job 237 (32800)
- job 231 (60000)
- job 246 (27000)
Total work in progress at the end of August $376050

3. Following is the direct material cost and applied overhead for job 248 :-

Direct material cost (see calculation in note3 below ) $331165
Applied overhead (see calculation in note 2 below ) $21385

Notes :-

Calculation of applied overhead

We know applied overhead rate = $0.91 per direct labor dollar

Direct labor in job 248 = $23500

Thus, applied overhead = 0.91x23500 = $21385.

Thus Thus direct material cost ,= Work in progress - direct labor - applied overhead

= 376050 (from note 1)-23500-21385 = $331165.